OpenStudy (anonymous):

Solve the given equation sin^2 θ = 6 − 2 cos^2 θ

OpenStudy (anonymous):

Use the Pythagorean trigonometric identity

OpenStudy (anonymous):

Sin^2(x)+Cos^2(x)=6-Cos^2(x) 1=6-Cos^2(x)

OpenStudy (anonymous):

I have it all the way down to sinθ=sqrt2/2

OpenStudy (anonymous):

2sin^2θ+sin^2θ-6=0 3sin^2θ-6 sin^2θ=2 sinθ=sqrt2/2 am I wrong?

OpenStudy (anonymous):

I'm not completely sure, I get a complex result trying to solve it, hehe

OpenStudy (anonymous):

can you show me what you have?

OpenStudy (anonymous):

sin^2(x)=6-2cos^2(x) sin^2(x)=6-2*(1-sin^2(x)) -4=sin^2(x)

OpenStudy (anonymous):

sinx=2i

OpenStudy (turingtest):

noliec that is not how you factor...

OpenStudy (anonymous):

Please correct me then, I'm only trying to get better and help people.

OpenStudy (turingtest):

actually it's not the factoring where you messed up, I apologize I misread your solution let me try to find the problem here

OpenStudy (turingtest):

sorry, I'm getting the same issue as noliec, which suggests no real solutions

OpenStudy (anonymous):

No problems TuringTest, thanks for confirming my thoughts!

OpenStudy (anonymous):

can you show me how did you get to no solution it may help with future problems

OpenStudy (turingtest):

yes http://www.wolframalpha.com/input/?i=sin%5E2x%3D6-2cos%5E2x&t=crmtb01 no real solutions -again, I apologize to Noliec I got confused and thought you factored out a 2 when yuo shouldn't have

OpenStudy (turingtest):

this can be shown to have no real solution in the way noliec did

OpenStudy (anonymous):

sin^2(x)=-4 Sin(x)=2i x=sinh^-1(2)*i

OpenStudy (anonymous):

OMG i was looking at that earlier and it didnt even click in my head

OpenStudy (turingtest):

also you can reduce this to \[\sin^2x=-4\implies\frac12(1-\cos2\theta)=-4\implies\cos(2\theta)=9\]and cos cannot be greater than 1, so there is no solution